Susan can paint a house in 4 hours. James can paint the same house in 6 hours. How long will it take to paint the house if Susan and James work together?





Found 2 solutions by Aust05, Alan3354:
Answer by Aust05(12)

Around 2 hours 20 min do this pretend sus an paints 1 pph or one paintstroke per hour so in 4 hours she does 4 paintstrokes and to paint the house you need to do four paintstroke james does 2/3 pph so that for every hour for every 20 min susan does 3/9of the stroke and james does 2/9 and the total painstroke needed is 36/9 by adding every twenty minutes you find u add 7 times or 2 hrs twenty mins
3/9*7=21/9
2/9*7=14/9
14/9+21/9=35/9 you want 36/9 so this is close enough good luck

Answer by Alan3354(69443)

Susan can paint a house in 4 hours. James can paint the same house in 6 hours. How long will it take to paint the house if Susan and James work together?
---------------------
Susan does 1/4 house per hour.
James does 1/6 house per hour.
Together, they do 1/4 + 1/6 per hour = 5/12
5/12 house/hour --> 12/5 hours/house
= 2.4 hours.
-----------------
Use the shortcut, product/sum
4*6/(4+6) = 24/10
= 2.4 hours